TheKayla:

Is this correct? 4x+3=8 Step 1: Subtract 3 from both sides. 4x+3−3=8−3 4x=5 Step 2: Divide both sides by 4. 4x 4 = 5 4 x= 5 4

Batman567:

No

TheKayla:

I believe it is the answer of 5/4

TheKayla:

@batman567 wrote:
No
Nobody's asking you newbie

montes67:

@thekayla wrote:
I believe it is the answer of 5/4
Thats what I got.

MidniteSunny:

Hmn well first of all, dis ish supposed to b posted in Mathematics, not Psychology also Step 2 ish a bit difficult for me to read and understand but I'll try anyways

TheKayla:

@midnitesunny wrote:
Hmn well first of all, dis ish supposed to b posted in Mathematics, not Psychology also Step 2 ish a bit difficult for me to read and understand but I'll try anyways
Idc it wouldn't let me

MidniteSunny:

Step 1, correct, Step 2 also correct if dat says chu got 5/4, or 1.25

freshavacado850:

are you solving for x or?

TheKayla:

Yes

TheKayla:

I need a second opinion. Not just a guarantee.

TheKayla:

Can you guys show work?

TheKayla:

I need to see it worked out

MidniteSunny:

4x+3=8 4x+3-3=8-3 4x=5 4x/4=5/4 x = 5/4 or 1.25

TheKayla:

@midnitesunny wrote:
4x+3=8 4x+3-3=8-3 4x=5 4x/4=5/4 x = 5/4 or 1.25
Thanks